[figure: 9.7 cm]

CRYSTAL-FIELD ORIGIN OF THE GIANT MAGNETOCRYSTALLINE ANISOTROPY OF PrRu_{2}Si_{2}

R. Michalski^{a,b}, Z. Ropka^{b} and R.J. Radwański^{a,b} aInst. of Physics, Pedagogical University, 30-084 Kraków bCenter for Solid State Physics, św. Filip 5, 31-150 Kraków, Poland.

PrRu_{2}Si_{2} is ferromagnetic below 14 K. It exhibits enormous magnetic anisotropy. A field of 5.5 T applied along the a-axis (the hard direction) induces only 1.3% of the moment observed along the easy axis (the tetragonal c-axis). It points to the anisotropy field of more than 400 T. We have performed calculations of the fine electronic structure of the Pr^{3+} ion in the tetragonal symmetry, relevant to PrRu_{2}Si_{2}, taking into account crystal-field and inter-site exchange interactions. Our calculations reproduce well the temperature dependence of the magnetic susceptibility and the heat capacity, in particular the λ-type of peak at T_{C}. We have calculated magnetization curves that yield so large anisotropy field as 400 T and the magnetocrystalline-anisotropy energy K_{1} of 59 J/cm^{3}. It is the largest known anisotropy - the anisotropy energy of the Nd_{2}Fe_{14}B supermagnet amounts to 12.5 J/cm^{3} only. Unfortunatelly, this giant anisotropy is confined to low temperatures only what prohibits its technical applications in the permanent-magnet industry. In conclusion, the giant anisotropy field as large as 400 T observed in PrRu_{2}Si_{2} can be accounted for by crystal-field interactions.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : oral

Corresponding author : R. Michalski

Address for correspondence : Center for Solid State Physics, św. Filip 5, 31-150 Kraków

Email address : sframich@wsp.krakow.pl